OpenStudy (anonymous):

Choose the system of equations which matches the following graph. MULTIPLE CHOICE

OpenStudy (anonymous):

5x + 2y = –14 7x + y = 7 5x + 2y = –14 7x – y = 7 5x – 2y = –14 7x + y = 7 5x – 2y = –14 7x – y = 7

OpenStudy (anonymous):

i would change all the equations to the standard slope form y = mx + b (where m is the slope and b is the y intercept) both the lines cut y-axis at -7 so look for the one that have y intercept = -7

OpenStudy (anonymous):

y = 5/2x - 7 y = 7x + 7

OpenStudy (anonymous):

y = 5/2x - 7 y = -7x - 7

OpenStudy (anonymous):

y = -5/2x + 7 y = 7x + 7

OpenStudy (anonymous):

for the first one it shuold be y = -5x/2 - 7 y = -7x + 7

OpenStudy (anonymous):

y = -5/2x + 7 y = -7x - 7

OpenStudy (anonymous):

the 2nd one should be y = -5x/2 - 7 y = 7x -7 you got the signs all wrong o.O

OpenStudy (anonymous):

3rd one: y = 5x/2 + 7 y = 7x + 7

OpenStudy (anonymous):

mannn

OpenStudy (anonymous):

haha, anyway 4th one y = 5x/2 + 7 y = 7x - 7 now look at all 4, see which one have y-intercept = -7 (hint: y = mx + b, b is the y-intercept)

OpenStudy (anonymous):

i believe the answer is B

OpenStudy (anonymous):

because both equations have a y intercept of -7 and on the graph the two lines intersect at 0,7

OpenStudy (anonymous):

ya. intersect at (0, -7) be careful with the signs next time :P b is correct

OpenStudy (anonymous):

thank you!